# Quantum theory of surface lattice resonances

### v1

@@ V Acknowledgments
^$\hat{\mathcal{H}}_{\text{int}}=\hbar\sum_{j,\nu<\nu^{\prime}}\sum_{\mathbf{k},\lambda}g_{\mathbf{k},\lambda}^{\nu\nu^{\prime}}\left(\hat{a}_{\mathbf{k},\lambda}^{\dagger}\mathrm{e}^{-\mathrm{i}\mathbf{k}\cdot\mathbf{r}_{j}}\hat{\sigma}_{j}^{\nu\nu^{\prime}}+\mathrm{H.c.}\right),$ (S47)
where $\hat{\sigma}_{j}^{\nu\nu^{\prime}}=\ket{\nu}_{j}\bra{\nu^{\prime}}_{j}$ describes the lowering operator between transitions $\nu$ and $\nu^{\prime}$ for a given molecule $j$ and the coupling with the electromagnetic field is given by $g_{\mathbf{k},\lambda}^{\nu\nu^{\prime}}=-\mathrm{i}\mathcal{E}_{\mathbf{k}}(\boldsymbol{\varepsilon}_{\mathbf{k},\lambda}\cdot\boldsymbol{\mu}_{\nu\nu^{\prime}})/\hbar$.
One can now follow the same procedure as detailed in Sec. S1. Derivation of non-Markovian lattice response in eliminating the photon modes to derive an effective equation for the matter part in momentum space. The dipole operator of the $\nu\nu^{\prime}$-transition can be expressed as
$-\mathrm{i}\omega\hat{\sigma}_{\mathbf{q}}^{\nu\nu^{\prime}}(\omega)$ $=-\left(\frac{\Gamma_{\nu\nu^{\prime}}}{2}+\mathrm{i}\omega_{\nu\nu^{\prime}}\right)\hat{\sigma}_{\mathbf{q}}^{\nu\nu^{\prime}}(\omega)+\mathrm{i}\mathcal{S}_{\mathbf{q},\nu\nu^{\prime}}(\omega)\hat{\sigma}_{\mathbf{q}}^{\nu\nu^{\prime}}(\omega)p_{\mathrm{inv}}^{\nu\nu^{\prime}}-p_{\mathrm{inv}}^{\nu\nu^{\prime}}\hat{\sigma}_{\mathbf{q}}^{\nu\nu^{\prime},\mathrm{in}}(\omega),$ (S48)
@@
Importantly, here we have made the assumption of equal and constant population in all sublevels $p_{j,\mathrm{inv}}^{\nu\nu^{\prime}}=\expectationvalue{\sigma_{j,\nu\nu^{\prime}}^{z}}\equiv p_{\mathrm{inv}}^{\nu\nu^{\prime}}$. Furthermore, we have neglected interactions among the different subtransitions by assuming that the frequencies of the different electronic transitions are very off-resonant. The lattice sum of the $\nu\nu^{\prime}$ transition dipoles is given by
$\mathcal{S}_{\mathbf{q},\nu\nu^{\prime}}(\omega)=3\pi\Gamma_{\nu\nu^{\prime}}^{\mathrm{rad}}c\,\frac{\omega^{2}}{\omega_{\nu\nu^{\prime}}^{3}}\left(\sum_{\Delta=1}^{M-1}\mathrm{e}^{-\mathrm{i}\mathbf{q}\cdot\mathbf{r}_{\Delta}}\boldsymbol{\varepsilon}_{\boldsymbol{\mu}_{\nu\nu^{\prime}}}\cdot\mathbf{G}(\mathbf{r}_{\Delta},\omega)\cdot\boldsymbol{\varepsilon}_{\boldsymbol{\mu}_{\nu\nu^{\prime}}}\right).$ (S50)
@@
$\hat{\sigma}_{\mathbf{q}}^{\nu\nu^{\prime}}(\omega)=\frac{-p_{\mathrm{inv}}^{\nu\nu^{\prime}}\hat{\sigma}_{\mathbf{q}}^{\nu\nu^{\prime},\mathrm{in}}(\omega)}{\mathrm{i}\left(\omega_{\nu\nu^{\prime}}-\omega\right)+\Gamma_{\nu\nu^{\prime}}^{\mathrm{rad}}/2+\mathrm{i}p_{\mathrm{inv}}^{\nu\nu^{\prime}}\mathcal{S}_{\mathbf{q},\nu\nu^{\prime}}(\omega)}.$ (S51)

### v2

@@ VII Acknowledgments
$\hat{\mathcal{H}}_{\text{int}}=\hbar\sum_{j,\nu<\nu^{\prime}}\sum_{\mathbf{k},\lambda}g_{\mathbf{k},\lambda}^{\nu\nu^{\prime}}\left(\hat{a}_{\mathbf{k},\lambda}^{\dagger}\mathrm{e}^{-\mathrm{i}\mathbf{k}\cdot\mathbf{r}_{j}}\hat{\sigma}_{j}^{\nu\nu^{\prime}}+\mathrm{H.c.}\right),$ (S47)
[…]
One can now follow the same procedure as detailed in Sec. S1 in eliminating the photon modes to derive an effective equation for the matter part in momentum space. The dipole operator of the $\nu\nu^{\prime}$-transition can be expressed as
$-\mathrm{i}\omega\hat{\sigma}_{\mathbf{q}}^{\nu\nu^{\prime}}(\omega)$ $=-\left(\mathrm{i}\omega_{\nu\nu^{\prime}}+\frac{\Gamma_{\nu\nu^{\prime}}}{2}\right)\hat{\sigma}_{\mathbf{q}}^{\nu\nu^{\prime}}(\omega)+\mathrm{i}\mathcal{S}_{\mathbf{q}}^{\nu\nu^{\prime}}(\omega)\hat{\sigma}_{\mathbf{q}}^{\nu\nu^{\prime}}(\omega)p_{\mathrm{inv}}^{\nu\nu^{\prime}}-p_{\mathrm{inv}}^{\nu\nu^{\prime}}\hat{\sigma}_{\mathbf{q}}^{\nu\nu^{\prime},\mathrm{in}}(\omega),$ (S48)
@@
Importantly, here we have made the assumption of equal and constant population in all sublevels $p_{j,\mathrm{inv}}^{\nu\nu^{\prime}}=\expectationvalue{\hat{\sigma}_{j,\nu\nu^{\prime}}^{z}}\equiv p_{\mathrm{inv}}^{\nu\nu^{\prime}}$. Furthermore, we have neglected interactions among the different subtransitions by assuming that the frequencies of the different electronic transitions are very off-resonant. The lattice sum of the $\nu\nu^{\prime}$ transition dipoles is given by
$\mathcal{S}_{\mathbf{q}}^{\nu\nu^{\prime}}(\omega)=3\pi\Gamma_{\nu\nu^{\prime}}^{\mathrm{rad}}c\,\frac{\omega^{2}}{\omega_{\nu\nu^{\prime}}^{3}}\left(\sum_{\Lambda\setminus\{0\}}\mathrm{e}^{-\mathrm{i}\mathbf{q}\cdot\mathbf{r}_{\Lambda}}\boldsymbol{\varepsilon}_{\boldsymbol{\mu}_{\nu\nu^{\prime}}}\cdot\mathbf{G}(\mathbf{r}_{\Lambda},\omega)\cdot\boldsymbol{\varepsilon}_{\boldsymbol{\mu}_{\nu\nu^{\prime}}}\right).$ (S50)
@@
$\hat{\sigma}_{\mathbf{q}}^{\nu\nu^{\prime}}(\omega)=\frac{-p_{\mathrm{inv}}^{\nu\nu^{\prime}}\hat{\sigma}_{\mathbf{q}}^{\nu\nu^{\prime},\mathrm{in}}(\omega)}{\mathrm{i}\left(\omega_{\nu\nu^{\prime}}-\omega\right)+\Gamma_{\nu\nu^{\prime}}^{\mathrm{rad}}/2+\mathrm{i}p_{\mathrm{inv}}^{\nu\nu^{\prime}}\mathcal{S}_{\mathbf{q}}^{\nu\nu^{\prime}}(\omega)}.$ (S51)
This corresponds to the linear response result, as will be shown in the following section S5.
S5. Pump-probe spectroscopy of nonlinear SLR dynamics
In this section, we show how to explicitly include population dynamics into our treatment. Neglecting cross-coupling between different dipole transitions, the equations of motion for the coherences express in time domain as
$\dot{\hat{\sigma}}_{j}^{12}(t)$ $=-\mathrm{i}\omega_{12}\hat{\sigma}_{j}^{12}-\hat{\sigma}_{j}^{12,\mathrm{in}}(t)\hat{p}_{\mathrm{inv},j}^{12}(t)+\hat{p}_{\mathrm{inv},j}^{12}(t)\sum_{\mathbf{k},\lambda}|g_{\mathbf{k},\lambda}^{12}|^{2}\sum_{j^{\prime}}\int_{0}^{t}\mathrm{d}t^{\prime}\mathrm{e}^{-\mathrm{i}\omega_{\mathbf{k}}(t-t^{\prime})}\hat{\sigma}_{j^{\prime}}^{12}(t^{\prime})\mathrm{e}^{\mathrm{i}\mathbf{k}\cdot(\mathbf{r}_{j}-\mathbf{r}_{j^{\prime}})},$ (S52a)
$\dot{\hat{\sigma}}_{j}^{23}(t)$ $=-\mathrm{i}\omega_{23}\hat{\sigma}_{j}^{23}-\hat{\sigma}_{j}^{23,\mathrm{in}}(t)\hat{p}_{\mathrm{inv},j}^{23}(t)+\hat{p}_{\mathrm{inv},j}^{23}(t)\sum_{\mathbf{k},\lambda}|g_{\mathbf{k},\lambda}^{23}|^{2}\sum_{j^{\prime}}\int_{0}^{t}\mathrm{d}t^{\prime}\mathrm{e}^{-\mathrm{i}\omega_{\mathbf{k}}(t-t^{\prime})}\hat{\sigma}_{j^{\prime}}^{23}(t^{\prime})\mathrm{e}^{\mathrm{i}\mathbf{k}\cdot(\mathbf{r}_{j}-\mathbf{r}_{j^{\prime}})},$ (S52b)
[…]
$\partial_{t}[\hat{\sigma}_{j}^{12,\dagger}\hat{\sigma}_{j}^{12}](t)$ $=\hat{\sigma}_{j}^{12,\mathrm{in}}(t)\hat{\sigma}_{j}^{12,\dagger}(t)+\hat{\sigma}_{j}^{12,\mathrm{in}}(t)^{\dagger}\hat{\sigma}_{j}^{12}(t)-\left(\hat{\sigma}_{j}^{12,\dagger}(t)\sum_{\mathbf{k},\lambda}|g_{\mathbf{k},\lambda}^{12}|^{2}\sum_{j^{\prime}}\int_{0}^{t}\mathrm{d}t^{\prime}\mathrm{e}^{-\mathrm{i}\omega_{\mathbf{k}}(t-t^{\prime})}\hat{\sigma}_{j^{\prime}}^{12}(t^{\prime})\mathrm{e}^{\mathrm{i}\mathbf{k}\cdot(\mathbf{r}_{j}-\mathbf{r}_{j^{\prime}})}+\mathrm{H.c}\right)+\ldots,$ (S53)
where the $\ldots$ denotes other terms which become important beyond third order for our perturbative treatment below (third order coherences feeding back into populations) and which we therefore drop for the ease of notation.
Generally, this leads to an analytically intractable set of equations involving convolutions in the time domain and space. A simplification can be carried out by considering a perturbative expansion in terms of input amplitudes Mukamel (1995); Reitz et al. (2025). To this end, we consider input pulses
$\langle\hat{\sigma}_{j}^{12,\mathrm{in}}(t)\rangle$ $=\eta_{p}f_{p}(t)\,\mathrm{e}^{-\mathrm{i}\omega_{p}t}\mathrm{e}^{\mathrm{i}\mathbf{k}_{\parallel}\cdot\mathbf{r}_{j}},\qquad\langle\hat{\sigma}^{23,\mathrm{in}}_{j}(t)\rangle=\eta_{p^{\prime}}f_{p^{\prime}}(t-\tau_{\Delta})\,\mathrm{e}^{-\mathrm{i}\omega_{p^{\prime}}t}\mathrm{e}^{\mathrm{i}\mathbf{k}_{\parallel}\cdot\mathbf{r}_{j}},$ (S54)
where $\eta_{p,p^{\prime}}$ describes the pulse amplitudes (considered real here), $f_{p,p^{\prime}}(t)$ the temporal envelope (assumed Gaussian), $\tau_{\Delta}$ is the delay time between the pulses, and we assume collinear incidence of pump and probe pulses at wavevector $\mathbf{k}_{\parallel}$ parallel to the array plane. We then go to momentum space for both coherences and population operators
$\hat{\sigma}_{j}^{\nu\nu^{\prime}}=\frac{1}{\sqrt{M}}\sum_{\mathbf{q}}\hat{\sigma}_{\mathbf{q}}^{\nu\nu^{\prime}}\mathrm{e}^{\mathrm{i}\mathbf{q}\cdot\mathbf{r}_{j}},\quad[\hat{\sigma}^{\nu\nu^{\prime},\dagger}\hat{\sigma}^{\nu\nu^{\prime}}]_{j}=\frac{1}{\sqrt{M}}\sum_{\mathbf{q}}[\hat{\sigma}^{\nu\nu^{\prime},\dagger}\hat{\sigma}^{\nu\nu^{\prime}}]_{\mathbf{q}}\mathrm{e}^{\mathrm{i}\mathbf{q}\cdot\mathbf{r}_{j}},$ (S55)
and expand all operators in terms of their contribution in the input fields as
$\langle\hat{O}_{\mathbf{q}}\rangle=\sum_{n,m=0}^{\infty}\eta_{p}^{n}\eta_{p^{\prime}}^{m}O_{\mathbf{q}}^{(n)(m)}.$ (S56)
At each order, a closed set of equations for the expansion coefficients $O_{\mathbf{q}}^{(n)(m)}$ can then be derived. In particular, we will consider the modification of the probe response due to the second-order populations generated by the pump, corresponding to order $(2)(1)$ [see Feynman diagrams in Fig. S2].
First order in pump: $(1)(0)$.
[…]
$-\mathrm{i}\omega\sigma_{\mathbf{q}}^{12}(\omega)^{(1)(0)}$ $=-\left(\mathrm{i}\omega_{12}+\frac{\Gamma_{12}^{\mathrm{rad}}}{2}\right)\sigma_{\mathbf{q}}^{12}(\omega)^{(1)(0)}-\mathrm{i}\mathcal{S}_{\mathbf{q}}^{12}(\omega)\sigma_{\mathbf{q}}^{12}(\omega)^{(1)(0)}p_{\mathrm{inv}}^{12}(0)-\delta_{\mathbf{q},\mathbf{k}_{\parallel}}f_{p}(\omega-\omega_{p})p_{\mathrm{inv}}^{12}(0),$ (S57)
where the populations are fixed by their initial conditions $p_{\mathrm{inv}}^{12}(0)=p_{\mathrm{inv},j}^{12}(0)=[p_{\mathrm{inv},j}^{12}]^{(0)(0)}$, and $f_{p}(\omega)$ denotes the Fourier transform of $f_{p}(t)$. For initial condition $p_{\mathrm{inv},j}^{12}(0)=-1$ (all population initially in the ground state), we obtain
$-\mathrm{i}\omega\sigma_{\mathbf{q}}^{12}(\omega)^{(1)(0)}$ $=-\left(\mathrm{i}\omega_{12}+\frac{\Gamma_{12}^{\mathrm{rad}}}{2}\right)\sigma_{\mathbf{q}}^{12}(\omega)^{(1)(0)}+\mathrm{i}\mathcal{S}_{\mathbf{q}}^{12}(\omega)\sigma_{\mathbf{q}}^{12}(\omega)^{(1)(0)}+\delta_{\mathbf{q},\mathbf{k}_{\parallel}}f_{p}(\omega-\omega_{p}).$ (S58)
Therefore, at linear order only the momentum mode $\mathbf{q}=\mathbf{k}_{\ell,\parallel}$ is driven (and $\mathbf{q}=-\mathbf{k}_{\ell,\parallel}$ for the complex conjugate coherence $\sigma_{\mathbf{q}}^{12,*}$).
Second order in pump: $(2)(0)$.
[…]
$-\mathrm{i}\omega\left[\sigma^{12,\dagger}\sigma^{12}\right]_{\mathbf{q}}(\omega)^{(2)(0)}$ $=-2\mathrm{Im}\sum_{\mathbf{p}}\left[\sigma_{\mathbf{p}}^{12,*}(\omega)^{(1)(0)}\ast\bar{\mathcal{S}}_{\mathbf{q}-\mathbf{p}}^{12}\sigma_{\mathbf{q}-\mathbf{p}}^{12}(\omega)^{(1)(0)}\right]$ (S59)
$+\left[\sigma_{\mathbf{q}-\mathbf{k}_{\parallel}}^{12,*}(\omega)^{(1)(0)}\ast f(\omega-\omega_{p})\right]+\left[\sigma_{\mathbf{q}+\mathbf{k}_{\parallel}}^{12}(\omega)^{(1)(0)}\ast f(\omega-\omega_{p})\right],$
generally leading to convolutions in both momentum and frequency space. Here, an asterisk in the exponent describes complex conjugation and an asterisk between two symbols denotes convolution. Making use of the fact that the only momentum mode driven for the coherence at linear order is $\mathbf{k}_{\parallel}$ (and the opposite sign for the complex conjugate), we can see that only the zero-momentum mode is driven for the population at second order (see Fig. S2)
$-\mathrm{i}\omega\left[\sigma^{12,\dagger}\sigma^{12}\right]_{\mathbf{q}=0}(\omega)^{(2)(0)}$ $=-2\mathrm{Im}\left[\sigma_{-\mathbf{k}_{\parallel}}^{12,*}(\omega)^{(1)(0)}\ast\bar{\mathcal{S}}_{\mathbf{k}_{\parallel}}^{12}\sigma_{\mathbf{k}_{\parallel}}^{12}(\omega)^{(1)(0)}\right]$ (S60)
$+\left[\sigma_{-\mathbf{k}_{\parallel}}^{12,*}(\omega)^{(1)(0)}\ast f(\omega-\omega_{p})\right]+\left[\sigma_{\mathbf{k}_{\parallel}}^{12}(\omega)^{(1)(0)}\ast f(\omega-\omega_{p})\right].$
In analogy to Sec. S1, we pull the self-interaction term at $\mathbf{r}_{\Lambda}=0$ out of the lattice sum $\bar{\mathcal{S}}_{\mathbf{q}}^{12}$ [see Eq. (S23)], giving rise to radiative damping for the population:
$-2\mathrm{Im}\left[\left(\frac{\mathrm{i}\Gamma_{\mathrm{rad}}^{12}}{2}\right)\left(\sigma_{-\mathbf{k}_{\parallel}}^{12,\ast}(\omega)^{(1)(0)}\ast\sigma_{\mathbf{k}_{\parallel}}^{12}(\omega)^{(1)(0)}\right)\right]$ $=-\Gamma_{\mathrm{12}}^{\mathrm{rad}}\mathrm{Re}\left[\sigma_{-\mathbf{k}_{\parallel}}^{12,\ast}(\omega)^{(1)(0)}\ast\sigma_{\mathbf{k}_{\parallel}}^{12}(\omega)^{(1)(0)}\right]$
$-\Gamma_{\mathrm{12}}^{\mathrm{rad}}\left[\sigma_{-\mathbf{k}_{\parallel}}^{12,\ast}(\omega)^{(1)(0)}\ast\sigma_{\mathbf{k}_{\parallel}}^{12}(\omega)^{(1)(0)}\right],$ (S61)
where the last step holds since $\sigma^{*}\ast\sigma$ is the Fourier transform of a real time‑domain quantity and hence has Hermitian symmetry.
Therefore, in total we obtain for the second order populations
$-\mathrm{i}\omega\left[\sigma^{12,\dagger}\sigma^{12}\right]_{\mathbf{q}=0}(\omega)^{(2)(0)}$ $=-\Gamma_{\mathrm{12}}^{\mathrm{rad}}\left[\sigma_{-\mathbf{k}_{\parallel}}^{12,\ast}(\omega)^{(1)(0)}\ast\sigma_{\mathbf{k}_{\parallel}}^{12}(\omega)^{(1)(0)}\right]-2\mathrm{Im}\left[\sigma_{-\mathbf{k}_{\parallel}}^{12,*}(\omega)^{(1)(0)}\ast\bar{\mathcal{S}}_{\mathbf{k}_{\parallel}}^{12}\sigma_{\mathbf{k}_{\parallel}}^{12}(\omega)^{(1)(0)}\right]$ (S62)
$+\left[\sigma_{-\mathbf{k}_{\parallel}}^{12,*}(\omega)^{(1)(0)}\ast f(\omega-\omega_{p})\right]+\left[\sigma_{\mathbf{k}_{\parallel}}^{12}(\omega)^{(1)(0)}\ast f(\omega-\omega_{p})\right].$
Second order in pump, first order in probe: $(2)(1)$.
These second-order populations can now act as the generator of third-order coherences $\sigma_{\mathbf{q}}^{23}(\omega)^{(2)(1)}$ (by creating a non-zero inversion in the $2\!\leftrightarrow\!3$ transition) via application of the probe field $p^{\prime}$:
$-\mathrm{i}\omega\sigma_{\mathbf{q}}^{23}(\omega)^{(2)(1)}$ $=-\left(\mathrm{i}\omega_{23}+\frac{\Gamma_{23}^{\mathrm{rad}}}{2}\right)\sigma_{\mathbf{q}}^{12}(\omega)^{(2)(1)}+\left[f_{p^{\prime}}(\omega-\omega_{p^{\prime}})\ast\left(\sigma^{12,\dagger}\sigma^{12}\right)_{\mathbf{q}-\mathbf{k}_{\parallel}}(\omega)^{(2)(0)}\right](\omega)+\mathrm{i}\mathcal{S}_{\mathbf{q}}^{23}(\omega)\sigma_{\mathbf{q}}^{23}(\omega)^{(2)(1)}.$ (S63)
Therefore, the momentum mode that is driven for the coherence at third order is $\mathbf{q}=\mathbf{k}_{\parallel}$, since only the population with zero quasi-momentum is contributing.
[…]